OpenStudy (mony01):

Suppose that 2 J of work is needed to stretch a spring from its natural length of 32 cm to a length of 46 cm. (a) How much work is needed to stretch the spring from 40 cm to 42 cm?

ganeshie8 (ganeshie8):

\(\large F = -kx\) \(\large W = \int_0^x -kx dx = \frac{kx^2}{2} \)

ganeshie8 (ganeshie8):

substutute x = 46-32 = 14, and find the value of k

ganeshie8 (ganeshie8):

W = 2J, when x = 14 : 2 = k*14^2/2 k = ?

OpenStudy (mony01):

is k=2/98?

ganeshie8 (ganeshie8):

yes, u may simplify further by canceling 2 if u want.. but it doesnt matter

ganeshie8 (ganeshie8):

so the required equation for Work doen is : W = kx^2/2 W = 2/98 * x^2/2 W = x^2/98

ganeshie8 (ganeshie8):

(a) How much work is needed to stretch the spring from 40 cm to 42 cm? plugin x = 42-40 = 2

ganeshie8 (ganeshie8):

W = 2^2/98 = 4/98 J

OpenStudy (mony01):

so the answer is 4/98?

ganeshie8 (ganeshie8):

yup ! simplify a bit canceling 2 if u want..

OpenStudy (mony01):

it says its wrong